# Is there a dynamical tendency in $H_{0}$ with late time measurements?

### v1

@@ 1 Introduction
Our current understanding of the Universe relies on the $\Lambda$CDM model which incorporates the cosmological constant $\Lambda$ and a cold dark matter component. This model is the most accepted and successful in explaining most cosmological observations [1, 2]. One of the most important challenges within the astrophysics and cosmology community is the discrepancy of more than $4\sigma$ for the Hubble constant $H_{0}$, between local measurements using the cosmic distance ladder and early-time measurements of the cosmic microwave background (CMB). The former based on direct observations of Supernovae Type Ia (SNe Ia) calibrated with Cepheids by the SH0ES team [3] and the latter obtained by the Planck satellite, which relies on the $\Lambda$CDM model [1]. This so called Hubble tension has led to an exhaustive search for potential explanations. A wide range of models have been used to address the $H_{0}$ tension, proposing either a dark energy (DE) component during the early evolution of the Universe [4, 5, 6, 7, 8, 9], a DE component with a time-varying equation of state [10, 11, 12], extra interactions between the components of the Universe [13, 14, 15, 16], or modified gravity [17, 18, 19]. For a more comprehensive study about the $H_{0}$ tension and alternative proposals, see the reviews [20, 21, 22].
@@
Recently, the Dark Energy Spectroscopic Instrument (DESI) collaboration published its cosmological constraints from the first year of observations of baryon acoustic oscillations (BAO). The results favor models with a DE component whose equation of state varies with time by more than $2.6\sigma$ [35], which challenges the paradigm of what we previously believed to be correct in the $\Lambda$CDM model. This dynamical DE could be responsible for a $H_{0}$ that depends on redshift.
Ref. [25] found that splitting cosmological data in different redshift bins up to $z=0.7$ gives different estimates to $H_{0}$ which are in contradiction with a single constant $H_{0}$ at $2.1\sigma$ level. Leading to the conclusion that the Hubble tension is present even between late time measurements themselves when taken at different redshifts. Ref. [25] assumed a cosmological constant and used BAO-BOSS data [2]. The new DESI results lead us to question whether the new data together with a dynamical DE could solve this tension.
The paper is organized as follows. In section 2, we detail the implemented methodology, along with the theory behind the cosmological probes used in this paper. In section 3, we mention the data used. In section 4, we explain the criteria used to split the data into bins. In section 5, we present our results on constraining the value of $H_{0}$ in each of the bins. Finally, the summary and conclusions can be found in section 6.
@@ 2 Cosmology
Following [25], we will split different cosmological data according to their redshift in order to estimate the Hubble constant using only data in certain redshift ranges. This method involves binning a modern dataset that includes the new DESI data, cosmic chronometers (CC), megamasers, and SNe Ia samples. The SNe Ia samples comprises the Pantheon+ compilation, and the new Union3 compilation and Dark Energy Survey (DES) Year 5 data release. With this new data we were able to expand the redshift range in the analysis to observe in more depth the evolution of $H_{0}$ with the redshift of the data.
@@
The water megamasers found in the accretion disks of supermassive black holes in active galactic nuclei (AGN) act like a laser beam in the microwave band, providing a method to calculate extragalactic distances without relying on the cosmic distance ladder [39]. The disk model returns an estimation of the angular diameter distance $\hat{D}_{\text{A}}$ to the galaxy and its recession velocity $\hat{v}_{i}$. The decomposition of this velocity between its peculiar and cosmological components $\hat{v}_{i}=v_{\textrm{pec,i}}+cz_{i}$ is unknown for objects this close. Each $\hat{D}_{\text{A}}$ is related to its expected redshift $z_{i}$ and the cosmological parameters by the equation
@@
Following [40], we take peculiar velocities of the galaxies into account incorporating $\sigma_{\textrm{pec}}=250$ km $\textrm{s}^{-1}$ into the velocity uncertainties. The chi-squared used to constrain cosmological parameters is then given as
@@
The SNe Ia data gives the apparent magnitude $m_{\textrm{b}}$ of each SNe, which can be compared with the predicted model given by
@@
$z_{\textrm{hel}}$ is the heliocentric redshift of the SNe (or its host galaxy) and $z_{\textrm{cmb}}$ is this redshift corrected for the peculiar motion of the solar system with respect to the CMB. Note that the parameters M and $H_{0}$ are degenerate when analyzing SNe alone. Therefore, SNe measurements are used to constrain other cosmological parameters, particularly the matter density $\Omega_{m}$.
@@
These measurements can also be summarized by a single estimation representing the spherically averaged distance
@@
BAO measurements constrain these distances all divided by the sound horizon $r_{\textrm{d}}$. Therefore, $H_{0}$ and $r_{\textrm{d}}$ are degenerated with each other. The BAO measurements constraint the product $H_{0}r_{\textrm{d}}$, and therefore are key in the treatment of Hubble tension. This constraint implies that a higher $H_{0}$, as that obtained by SH0ES [3], requires a lower $r_{\text{d}}$, which is incompatible with the value inferred by the Planck Collaboration [1]. In this work we won’t use the CMB prior on $r_{\textrm{d}}$ as this will imply the use of data outside our redshift bins.
@@ 3 Data
    For the superenovae type Ia measurements, we use any of the three latest compilations. 1,590 SNe Ia from the Pantheon+ compilation in the redshift range $0.01\leq z\leq 2.26$ [53]. Recently published results from the Year 5 data release of the Dark Energy Survey (DES Y5) that include 1,635 new photometrically classified SNe Ia and 194 low-redshift SNe Ia (which are also in common with Pantheon+), spanning $0.025\leq z\leq 1.3$ [54]. Finally, 22 binned data points from 2087 SNe Ia from the Union3 sample in the range $0.05\leq z\leq 2.26$ [55], which has 1360 SNe in common with Pantheon+. We include both the statistical and systematic uncertainties in the three sets. We emphasize that these SNe Ia data are not independent of each other, and therefore, we cannot combine them. However, they differ in the way they analyze systematic errors and astrophysical parameters, making it interesting to compare the results of each of them.
Compared with the data used in ref. [25] which includes the BOSS BAO data [2] and the Pantheon sample [56], we have the BAO data from DESI Y1, the recent supernovae compilations Pantheon+, Union3 and DES Y5, as well as three extra measurements of CC.
Since the Pantheon+, DES and Union3 data cannot be combined, we present each of their results separately. The analyses always include the megamasers+CC+DESI data, which we will call “Base”, but differ in the SNe sample. We denote each case as Base+PantheonPlus, Base+Union3 and Base+DES.
@@ 4 Fitting
In this section, we describe how we fit the cosmological parameters using the data specified in the last section. Our objective is to see if the estimation of the Hubble constant $H_{0}$ is consistent when using data at different redshifts. We will split the total redshift range from our data in separate redshift bins with $z\in(z^{i}_{min},z^{i}_{max})$ and we use only data coming from that bin to estimate $H_{0}$.
The redshift width of the bins varies between each other as we follow the strategy used in [25] requiring that the average redshift of each bin coincides between megamasers, SNe and BAO. We use the weighted average $\bar{z}$ given by
@@
where $N_{i}$ denotes the number of data used in the bin $i$ and $\sigma_{k}$ denotes the error in the observable at redshift $z_{k}$. For megamasers we use $\sigma_{k}^{2}=\sigma_{\hat{v},k}^{2}+\sigma_{\textrm{pec}}^{2}+\sigma_{\hat{D},k}^{2}$. The bins constructed in each analysis are summarized in table 1.
Not all bins contain all the data. The megamasers appear only in the first redshift bin, while some bins lack DESI data. We only require that all bins contain either CCs or megamasers to break the degeneracies needed to obtain a value of $H_{0}$. We emphasize that the condition (4.1) leads to different binnings and different $\bar{z}$’s depending on the SNe’s used (PantheonPlus, Union3 and DES).
@@
On the other hand, the data contained in a single bin lack the power to constrain the equation of state parameters $w_{0}$ and $w_{a}$. Therefore the posteriors of these parameters are mainly influenced by their priors. We considered large prior ranges
@@
Not all bins fit all parameters due to the fact that each bin may or may not contain a certain type of data. The only parameters fitted in all bins are $\Omega_{m}$, $H_{0}$, $w_{0}$, $w_{a}$, and the absolute magnitude of SNe Ia $M$.
@@ 5 Results
We show the fitted values of $H_{0}$ for each bin in Table 2 and illustrate them in figure 1. In all three analyses, $H_{0}$ decreases as function of redshift for $z<0.6$, which agrees with [25]. In the next two bins, the value of $H_{0}$ increases to a maximum of $\sim 78$ km $\text{s}^{-1}$ $\text{Mpc}^{-1}$ and then decreases again. Note that the scarcity and low quality of data inflate the errors at high redshifts.
@@
We evaluate how statistically significant is the deviation from a constant $H_{0}$ by using three different methods. First, we use the Bayesian evidence given by [58]
@@
where $\Pi$ is the prior and $\mathscr{L}$ the Likelihood of the parameters $\theta_{k}$. We compute the integral using the nested integration algorithm Polychord [59, 60]. Here a larger value means a model with more probability. We report the Bayes factor given by
@@
where $\mathscr{L}_{\text{max}}$ is the maximum likelihood that can be obtained within the model and $k$ the number of free parameters. Between the two models, the variable and the constant, the one with the smaller AIC value is preferred by the data. This criterion favors the larger $\mathscr{L}_{\text{max}}$ but penalizes too many extra parameters. In the next subsections we report $\Delta\text{AIC}=\text{AIC}_{\text{var}}-\text{AIC}_{\text{cons}}$.
Thirdly, we determine the number of standard deviations ($\sigma$) away from a constant. For parameterizations with a single extra parameter we only divide the distance to a constant by the standard deviation of the parameter. For models with more parameters, we compute the $p$-value of the variable model and, assuming a Gaussian distribution, we determine the number of standard deviations associated with that $p$-value. This is the least rigorous of our tests, but serves to communicate the relevance of our results as the cosmological community is familiar with the concept that a few $\sigma$’s can be the result of statistical fluctuations and not a physical effect.
@@ 5.1 Quadratic parameterization
From the table 3 we can see that the parameters $m$ and $b$ deviate from zero, favouring a non constant $H_{0}$ model. The values of the Bayes factor and the $\Delta$AIC however, show that this parametrization improves very modestly the fit to the data given by the constant $H_{0}$. Only accounting for a barely worth mentioning result according to the Jeffreys’ scale [63] of the Bayes factor. And according to the $\Delta$AIC scale in [62] corresponding to no preference between the models.
It is evident that the low values of $H_{0}$ in the last two bins can lead to a fitting which is compatible with a constant $H_{0}$ ($m=0$ and $b=0$). This can be considered a limitation of the quadratic parameterization, which asks for $H_{0}$ to grow up to infinity at large redshifts. As this parameterization is purely phenomenological, we change it for one that won’t present this behaviour in the next subsection.
@@ 5.2 Fourier parameterization
We now fit all the available bins showing the results in table 4. Figure LABEL:fig:ajustes_modelo_cos2x displays the fit for the 8 bins in the Base+PantheonPlus and Base+Union3 samples, and for only 6 bins in the Base+DES sample, remembering this data set has a reduced redshift range.
Again, table 4 includes the $\sigma$, $\Delta$AIC and $B$ values to see the robustness of this model against a constant $H_{0}$. The parameter $A$ is different from zero up to $\sim 2.3$ standard deviations, suggesting a preference for the non constant model. However, the scale in [62] states that the difference in AIC values is insufficient to favor any of the models. Table 4 also contains the results in each data set for only 6 bins, confirming a general improvement over the quadratic parameterization. On Jeffreys’ scale [63], the Bayes factors in table 4 account for a barely worth mentioning value, with the exception of the Base+DES sample, where we found a substantial preference for the non constant model. It is interesting to note that the latest two bins in Base+PantheonPlus and Base+Union3 don’t change too much the results in each data set compared to 6 bins.
@@ 5.3 Unbinned results
Data ${H}_{0}\left[\frac{\text{km}}{\text{s Mpc}}\right]$ $\Omega_{m}$ $w_{o}$ $w_{a}$ $r_{d}$ [Mpc]

### v2

@@ 1 Introduction
Our current understanding of the Universe relies on the $\Lambda$CDM model, which incorporates the cosmological constant $\Lambda$ and a cold dark matter component. This model is the most accepted and successful in explaining most cosmological observations [1, 2]. One of the most important challenges within the astrophysics and cosmology community is the discrepancy of more than $4\sigma$ for the Hubble constant $H_{0}$, between local measurements using the cosmic distance ladder and early-time measurements of the cosmic microwave background (CMB). The former was based on direct observations of Supernovae Type Ia (SNe Ia) calibrated with Cepheids by the SH0ES team [3] and the latter was obtained by the Planck satellite, which relies on the $\Lambda$CDM model [1]. This so-called Hubble tension has led to an exhaustive search for potential explanations. A wide range of models have been used to address the $H_{0}$ tension, proposing either a dark energy (DE) component during the early evolution of the Universe [4, 5, 6, 7, 8, 9], a DE component with a time-varying equation of state [10, 11, 12], extra interactions between the components of the Universe [13, 14, 15, 16], or modified gravity [17, 18, 19]. However, in most solutions, reducing the tension leads to a mismatch with other well-measured cosmological quantities or fails to fully address the Hubble tension, so no proposal so far presents solid arguments to claim it is better than all the others [20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. For a more comprehensive study about the $H_{0}$ tension and alternative proposals, see the reviews [30, 31, 32].
@@
Recently, the Dark Energy Spectroscopic Instrument (DESI) collaboration published its cosmological constraints from the first year of observations. The results favor models with a DE component whose equation of state varies with time by more than $2.6\sigma$ [45] for the baryon acoustic oscillations (BAO) measurements with similar results from the more recent full shape analysis [46] with a significance from $2.5\sigma$ up to $3.8\sigma$ depending on the Supernova data used. This challenges the paradigm of what we previously believed to be correct in the $\Lambda$CDM model. This dynamical DE could be responsible for a $H_{0}$ that depends on the redshift.
Ref. [35] found that splitting cosmological data in different redshift bins up to $z=0.7$ gives different estimates to $H_{0}$ which are in contradiction with a single constant at $2.1\sigma$ level. This leads to the conclusion that the Hubble tension is present even between late-time measurements themselves when taken at different redshifts. Ref. [35] assumed a cosmological constant and used BAO-BOSS data [2]. The new DESI results lead us to question whether the new data together with a dynamical DE could solve this tension.
Clearly, there is a need for model-independent techniques to reconstruct the expansion rate of the universe. Cosmography [47, 48, 49] is one of these approaches; since it only uses kinematic parameters together with the assumption of homogeneity and isotropy on large scales, it turns out to be a very useful tool to deal with degeneracies between cosmological models beyond $\Lambda$CDM and to understand the kinematics of our local universe. Letting the data constrain the Hubble rate without dealing with the problems related to the nature of dark energy or dark matter leads to an excellent consistency test to break the assumptions that were once established given the extremely limited data available for cosmology at that time. Then, after reconstructing $H(z)$ with the latest cosmological data we can ask if our cosmological models are consistent with it.
The paper is organized as follows. In Section 2, we detail the implemented methodology, along with the theory behind the cosmological probes used in this paper. In Section 3, we mention the data used. Section 4 explains the criteria used to split the data into bins. In Section 5, we present our results on constraining the value of $H_{0}$ in each of the bins. Finally, the summary and conclusions can be found in Section 6.
@@ 2 Cosmology
Following [35], we will split different cosmological data according to their redshift in order to estimate the Hubble constant using only data in certain redshift ranges. This method involves binning a modern dataset that includes the new DESI data, cosmic chronometers (CC), megamasers, and SNe Ia samples. The SNe Ia samples comprise the Pantheon+ compilation, the new Union3 compilation, and Dark Energy Survey (DES) Year 5 data release. With these new data, we were able to expand the redshift range in the analysis to observe in more depth the evolution of $H_{0}$ with the redshift of the data.
@@
The water megamasers found in the accretion disks of supermassive black holes in active galactic nuclei (AGN) act like a laser beam in the microwave band, providing a method to calculate extragalactic distances without relying on the cosmic distance ladder [53]. The disk model returns an estimate of the angular diameter distance $\hat{D}_{\text{A}}$ to the galaxy and its recession velocity $\hat{v}_{i}$. The decomposition of this velocity between its peculiar and cosmological components $\hat{v}_{i}=v_{\textrm{pec,i}}+cz_{i}$ is unknown for objects this close. Each $\hat{D}_{\text{A}}$ is related to its expected redshift $z_{i}$ and the cosmological parameters by the equation
@@
Following [54], we take peculiar velocities of the galaxies into account incorporating $\sigma_{\textrm{pec}}=250$ km $\textrm{s}^{-1}$ into the velocity uncertainties. The chi-squared used to constrain the cosmological parameters is then given as
@@
The SNe Ia data give the apparent magnitude $m_{\textrm{b}}$ of each SNe, which can be compared with the predicted model given by
@@
$z_{\textrm{hel}}$ is the heliocentric redshift of the SNe (or its host galaxy), and $z_{\textrm{cmb}}$ is this redshift corrected for the peculiar motion of the solar system with respect to the CMB. Note that the parameters M and $H_{0}$ are degenerate when analyzing SNe alone. Therefore, SNe measurements are used to constrain other cosmological parameters, particularly the matter density $\Omega_{m}$.
@@
These measurements can also be summarized by a single estimate representing the spherically averaged distance
@@
BAO measurements constrain these distances all divided by the sound horizon $r_{\textrm{d}}$. Therefore, $H_{0}$ and $r_{\textrm{d}}$ are degenerate with each other. BAO measurements constrain the product $H_{0}r_{\textrm{d}}$, and therefore are key in the treatment of Hubble tension. This constraint implies that a higher $H_{0}$, as that obtained by SH0ES [3], requires a lower $r_{\text{d}}$, which is incompatible with the value inferred by the Planck Collaboration [1]. In this work, we will not use the CMB prior on $r_{\textrm{d}}$ as this will imply the use of data outside our redshift bins.
@@ 3 Data
    For the superenovae type Ia measurements, we use any of the three latest compilations. 1,590 SNe Ia from the Pantheon+ compilation in the redshift range $0.01\leq z\leq 2.26$ [67]. Recently published results from the Year 5 data release of the Dark Energy Survey (DES Y5) that include 1,635 new photometrically classified SNe Ia and 194 low-redshift SNe Ia (which are also common to Pantheon+), spanning $0.025\leq z\leq 1.3$ [68]. Finally, 22 binned data points from 2087 SNe Ia from the Union3 sample in the range $0.05\leq z\leq 2.26$ [69], which has 1360 SNe in common with Pantheon+. We include both the statistical and systematic uncertainties in the three sets. We emphasize that these SNe Ia data are not independent of each other and therefore we cannot combine them. However, they differ in the way they analyze systematic errors and astrophysical parameters, making it interesting to compare the results of each of them.
Compared with the data used in ref. [35] which includes the BOSS BAO data [2] and the Pantheon sample [70], we have the BAO data from DESI Y1, the recent supernovae compilations Pantheon+, Union3, and DES Y5, as well as three extra measurements of CC.
Since the Pantheon+, DES, and Union3 data cannot be combined, we present each of their results separately. The analyzes always include the megamasers+CC+DESI data, which we will call “Base”, but differ in the SNe sample. We denote each case as Base+PantheonPlus, Base+Union3 and Base+DES.
@@ 4 Fitting
In this section, we describe how we fit the cosmological parameters using the data specified in the last section. The code used in this section is public at  [^2] See https://github.com/MauLoHdz/LUDB. Our objective is to see if the estimation of the Hubble constant $H_{0}$ is consistent when using data at different redshifts. We will split the total redshift range from our data into separate redshift bins with $z\in(z^{i}_{min},z^{i}_{max})$, and we use only data coming from that bin to estimate $H_{0}$.
The redshift widths of the bins vary between each other, as we follow the strategy used in [35] requiring that the average redshift of each bin coincides between megamasers, SNe and BAO. We use the weighted average $\bar{z}$ given by
@@
where $N_{i}$ denotes the number of data used in the bin $i$ and $\sigma_{k}$ denotes the error in the observable at redshift $z_{k}$. For megamasers, we use $\sigma_{k}^{2}=\sigma_{\hat{v},k}^{2}+\sigma_{\textrm{pec}}^{2}+\sigma_{\hat{D},k}^{2}$. The bins constructed in each analysis are summarized in table 1.
Not all bins contain all the data. The megamasers appear only in the first redshift bin, while some bins lack DESI data. We only require that all bins contain either CC or megamasers data to break the $H_{0}$-$r_{d}$ and $H_{0}$-M degeneracies, mentioned in section 2. We emphasize that the condition (4.1) leads to different binnings and different $\bar{z}$’s depending on the SNe’s used (PantheonPlus, Union3 and DES).
@@
On the other hand, the data contained in a single bin lack the power to constrain the equation of state parameters $w_{0}$ and $w_{a}$. Therefore, the posteriors of these parameters are mainly influenced by their priors. We considered large prior ranges
@@
Not all bins fit all parameters, since each bin may or may not contain a certain type of data. The only parameters fitted in all bins are $\Omega_{m}$, $H_{0}$, $w_{0}$, $w_{a}$, and the absolute magnitude of SNe Ia M.
@@ 5 Results
We show the fitted values of $H_{0}$ for each bin in Table 2 and illustrate them in Figure 1. In all three analyses, $H_{0}$ decreases as a function of the redshift for $z<0.6$, which is consistent with [35]. In the next two bins, the value of $H_{0}$ increases to a maximum of $\sim 78$ km $\text{s}^{-1}$ $\text{Mpc}^{-1}$ and then decreases again. Note that the scarcity and low quality of the data inflate the errors at high-redshifts.
@@
We evaluated how statistically significant the deviation from a constant $H_{0}$ is using three different methods. First, we use the Bayesian evidence given by [72]
@@
where $\Pi$ is the prior and $\mathscr{L}$ the Likelihood of the parameters $\theta_{k}$. We compute the integral using the nested integration algorithm, Polychord [73, 74]. Here, a larger value means a model with more probability. We report the Bayes factor given by
@@
where $\mathscr{L}_{\text{max}}$ is the maximum likelihood that can be obtained within the model and $k$ the number of free parameters. Between the two models, variable and constant, the one with the lowest AIC value is preferred by the data. This criterion favors the larger $\mathscr{L}_{\text{max}}$, but penalizes too many extra parameters. In the next subsections, we report $\Delta\text{AIC}=\text{AIC}_{\text{var}}-\text{AIC}_{\text{cons}}$.
Thirdly, we determine the number of standard deviations ($\sigma$) from a constant. For parameterizations with a single extra parameter, we only divide the distance to a constant by the standard deviation of the parameter. For models with more parameters, we compute the $p$-value of the variable model and, assuming a Gaussian distribution, we determine the number of standard deviations associated with that $p$-value. This is the least rigorous of our tests, but serves to communicate the relevance of our results as the cosmological community is familiar with the concept that a few $\sigma$’s can be the result of statistical fluctuations and not a physical effect.
@@ 5.1 Quadratic parameterization
From Table 3 we can see that the parameters $m$ and $b$ deviate from zero, favoring a non-constant $H_{0}$ model. However, the values of the Bayes factor and the $\Delta$AIC show that this parameterization very modestly improves the fit to the data given by the constant $H_{0}$. Only accounting for a barely worth mentioning result according to the Jeffrey’s scale [77] of the Bayes factor. And according to the $\Delta$AIC scale in [76], which corresponds to no preference between the models.
It is evident that the low values of $H_{0}$ in the last two bins can lead to a fitting that is compatible with a constant $H_{0}$ ($m=0$ and $b=0$). This can be considered a limitation of the quadratic parameterization, which asks for $H_{0}$ to grow up to infinity at large redshifts. As this parameterization is purely phenomenological, we replace it with one that will not exhibit this behavior in the next subsection.
@@ 5.2 Fourier parameterization
We now fit all the available bins showing the results in Table 4. Figure 3 displays the fit for the 8 bins in the Base+PantheonPlus and Base+Union3 samples, and for only 6 bins in the Base+DES sample, remembering that this dataset has a reduced redshift range.
Again, Table 4 includes the $\sigma$, $\Delta$AIC, and $B$ values to see the robustness of this model against a constant $H_{0}$. The parameter $A$ is different from zero up to $\sim 2.3$ standard deviations, suggesting a preference for the non-constant model. However, the scale in [76] indicates that the difference in the AIC values is insufficient to favor any of the models. Table 4 also contains the results in each dataset for only 6 bins, confirming a general improvement over the quadratic parameterization. On Jeffreys’ scale [77], the Bayes factor accounts for a barely worth mentioning value, with the exception of the Base+DES sample, where we found a substantial preference for the non-constant model. It is interesting to note that the latest two bins in Base+PantheonPlus and Base+Union3 don’t change too much the results in each dataset compared to 6 bins.
[…]
@@ 5.3 Unbinned results
Data ${H}_{0}\left[\frac{\text{km}}{\text{s Mpc}}\right]$ $\Omega_{m}$ $w_{0}$ $w_{a}$ $r_{d}$ [Mpc]
@@
5.4 $H(z)$ reconstruction
Finally, given this apparent variability in $H_{0}$, we wonder about deviations in the Hubble parameter, $H(z)$, from its theoretical prediction in the context of the CPL model. To do so, we will treat $H(z)$ as a free parameter to be constraint directly from the full data. Specifically, $H(z)$ will be determined by its values in a set of redshift nodes $H(z_{i})$, defined by equally spaced values of the scale factor $a_{i}$, starting at $a_{1}=1$. This choice is motivated by the fact that most of the data are in the range $z<1$, making it convenient for most nodes to be distributed within this region. The maximum redshift node will also be determined by the maximum of the previous bins in Table 1, i.e. $z_{\textrm{max}}=2.3$ for Base+PantheonPlus and Base+Union3, and $z_{\textrm{max}}=1.13$ for Base+DES, in the latter we used only the data with $z\leq 1.13$. We used 9 nodes for Base+PantheonPlus and Base+Union3, and 7 nodes for Base+DES.
Each of the nodes $H_{i}=H(z_{i})$ is treated as a free parameter, instead of the cosmological parameters $\Omega_{m}$, $\omega_{0}$ and $\omega_{a}$. The Hubble parameter at any $z$ interpolated with a cubic spline. Figure 5 shows the reconstruction and Table 6 contains the Hubble constant (first node), and compares the fitting of this model compared with the CPL model.
Base+DES has large error bars in its last node, due to the quality of data in this range. This has an immediate effect on the values of $\sigma$ and $\Delta$AIC, showing a preference for the CPL model. In contrast, using Base+PantheonPlus and Base+Union3, there is a preference for the reconstructed Hubble parameter over the CPL model, with Base+Union3 being the most notable, reaching a statistical significance of up to $3.5\sigma$ and a $\Delta$AIC value indicating considerable support according to the scale in [76] compared to the CPL model. In order to interpret the results, it is important to note the impact of the interpolation method used. Cubic splines can present overfitting in regions where the data is disperse, generating oscillations that could be interpreted as real features of the model, but are actually mathematical artifacts. However, the values of $\Delta$AIC were not significantly affected when we used quadratic or linear splines. This shows that the choice of the interpolation scheme does not have a significant effect on the reconstruction of the Hubble parameter in the range where the data lie.
Paying attention only to the inferred nodes $H(z_{i})$, we observe a slight tension around $z\sim 0.5$ for Base+PantheonPlus and Base+DES, whereas the others agree with its theoretical prediction.
Data $H_{0}\left[\frac{\text{km}}{\text{s Mpc}}\right]$ $\sigma$ $\Delta$AIC
Base+PantheonPlus $71.74^{+2.15}_{-2.30}$ 3.0 -1.91
Base+Union3 $73.28^{+2.53}_{-2.43}$ 3.5 -5.26
Base+DES $70.85^{+2.42}_{-2.24}$ 1.2 3.04
Table 6: The Hubble constant, $\sigma$ deviation, and $\Delta$AIC values between the reconstructed and the theoretical Hubble parameter. Base+DES favors the CPL model, probably due to the large error bars in the last node. For Base+PantheonPlus and Base+Union3, there is a preference for the reconstructed $H(z)$, as indicated by the $\sigma$ values, which rise up to $3.5\sigma$. The $\Delta$AIC values show no preference for either model from the scale in [76] for Base+PantheonPlus, but Base+Union3 shows considerable support against the theoretical predictions.
Figure 5: Reconstructed Hubble parameter $H(z)$ in each sample. The dashed line represents the interpolation obtained interpolating with a cubic spline the values of $H(z_{i})$ in each node, while the external black solid lines represent the $1\sigma$ and $2\sigma$ errors. The orange line shows the theoretical Hubble parameter following the CPL model, with the parameters values fixed from Table 5.